How do u solve 8.3+(-4.1)=?
Answer by Alan3354(69443)

How do u solve 8.3+(-4.1)=?
----------
It's not solving, just adding/subtracting.
--------
8.3+(-4.1) = 8.3 - 4.1
= 4.2